How do you DIVIDE this rational expression: 6-3x/5(division sign)4x-8/25? Whats the answer?: How do you DIVIDE this rational expression: 6-3x/5(division sign)4x-8/25? Whats the answer?
Answer by Fombitz(1755)

I'm assuminng you mean like this,
((6-3x)/5)/((4x-8)/25)
Dividing by a fraction is like multiplying by the reciprocal of the fraction.
So
(A/B)/(C/D)=(A/B)*(D/C)
So in your case,
((6-3x)/5)/((4x-8)/25)=((6-3x)/5)*(25/(4x-8))
((6-3x)/5)/((4x-8)/25)=((6-3x)/cross(5))*(cross(25)5/(4x-8))
((6-3x)/5)/((4x-8)/25)=(3(2-x))*(5/(4(x-2)))
((6-3x)/5)/((4x-8)/25)=(-3(x-2))*(5/(4(x-2)))
((6-3x)/5)/((4x-8)/25)=(-3*cross((x-2)))*(5/(4*cross((x-2))))
((6-3x)/5)/((4x-8)/25)=(-3)*(5/4)
((6-3x)/5)/((4x-8)/25)=-15/4